OpenStudy (anonymous):

How many cubic inches of ice cream can an ice cream cone hold (within the cone) if its height is 7 inches and the diameter of its base is 3 inches? Give solution accurate to the nearest hundredth. Include correct units with your solution.

OpenStudy (anonymous):

volume of a cone is 1/3 x area of base x height. area of base = pi x r^2. r = d/2. solve.

OpenStudy (anonymous):

k lemme me work it out

OpenStudy (anonymous):

You should also add that you're cheating yourself if you don't put a scoop on top of the cone.

OpenStudy (anonymous):

pls go away

OpenStudy (anonymous):

V=(1/3)π*r^2*h V=(1/3)π*(1.5)^2*(7) V=16.4934. The cone can hold 16.49 cubic inches of ice cream